Assume that you plan to take a housing loan with a tenor of 20 year. The loan has to be repaid in equal monthly installments. Considering that the loan amount is Rs. 50 lakhs and the interest rate on loan is 9% p.a., what would be the equated monthly installment (EMI)?


1
Expert's answer
2021-11-16T17:47:17-0500

In this we have to calculate present value factor and than can calculate the monthly payment.

Period =20 years

Total period"= 20 \times 12=240" months

Monthly interest rate"=\frac{9}{12}=0.75\%"


Present value factor

"=\frac{1−(1+r)^{−n}}{r}=\frac{1−1.0075^{−240}}{0.0075}=111.14495403"


Loan amount= Present value factor x monthly payment

5000000= 111.14495403 x monthly payment

Monthly payment=Rs 44986.30

would be the equated monthly installment (EMI).